OpenStudy (anonymous):

What is the prime factorization of 550?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (abmon98):

550 = 2 * 52 * 11 550 ÷ 2 = 275 275 ÷ 5 = 55 55 ÷ 5 = 11 11 ÷ 11 = 1

OpenStudy (abmon98):

i am sorry 5*5*2*11=550

OpenStudy (skullpatrol):

Prime factors of positive integers can be found by using the primes in order as divisors. The prime factorization of positive integer is the expression of the integer as a product of prime factors.

OpenStudy (anonymous):

550/5=110 110/11=10 10/2=5 5/5=1

OpenStudy (solomonzelman):

How about just keep dividing the number by whatever it seems divisible by, till you get a prime number ?

OpenStudy (skullpatrol):

But do it in order :-)

OpenStudy (solomonzelman):

Yes :)

OpenStudy (camerondoherty):

Try a factor tree

OpenStudy (solomonzelman):

I am going to show something:) 1) Click and hold ALT 2) click the number code (using the numbers that are on the right of the keyboard, and `NOT` the ones below `F1`, `F2`, `F3`, etc., ) 3) release the ALT number code result `0 2 1 5 ` × `2 4 6 ` ÷ `7 5 4 ` ≥ `7 5 5` ≤ `2 4 1 ` or `7 5 3` ± `2 4 7` ≈

OpenStudy (solomonzelman):

You can type any symbol you need without copy-pasting.

OpenStudy (midhun.madhu1987):